OpenStudy (anonymous):

8. To fence a rectangular farm of 750 m², 110 m of fence has been used. Calculate the dimensions of the farm. Please give me the steps also! Thanks! :D

OpenStudy (anonymous):

\[2\times (x+y) = 110\] \[x \times y = 750\] now solve the equations for x and y (the dimensions of the fence).

OpenStudy (anonymous):

More hints: -> x + y = 55 => y = 55 -x Thus x ( 55 -x ) = 750 Solve for x first!